JoelTheBoss (joel_the_boss):

(7 − 3i) • (2 − i) 4 + i 14 + 3i 11 − 13i 14 − 10i

JoelTheBoss (joel_the_boss):

@lolpizza.

OpenStudy (anonymous):

c

JoelTheBoss (joel_the_boss):

(−3 + 2i) • (2 + i) −8 + i −6 + 3i −3 + 4i −1 + 2i @lolpizza.

OpenStudy (anonymous):

I cant see what that is bc of the those question mark things

JoelTheBoss (joel_the_boss):

refresh

OpenStudy (anonymous):

ok got it, You can't solve it because it's not an equation. The best you can do is reduce it using FOIL: 3*2=6 3* -i = -3i 2i*2=4i 2i*-i= -2i^2 and i^2=-1 so this equals 2 (-2 * -1). Because it's not in parentheses, only the i gets squared, not the -2. Combine like terms to get 8+i

JoelTheBoss (joel_the_boss):

thanks :3

OpenStudy (anonymous):

youre welcome! anything else? :)

JoelTheBoss (joel_the_boss):

Simplify the given expression: (3 + 4i) + (5 − 2i) −2 + 6i 2 − 2i 7 + 3i 8 + 2i @lolpizza.

OpenStudy (anonymous):

(3 + (4 * i)) + (5 - (2 * i)) = 8 + 2 i

JoelTheBoss (joel_the_boss):

<33 omg your awesome lol

OpenStudy (anonymous):

ahhaa ty